NEGATIVE ION PHOTOELECTRON SPECTRA OF DIHALOCARBENES AND DIHALOMETHYL RADI-CALS

\underline{SCOTT WREN,} KRISTEN M. VOGELHUBER, and W. CARL LINEBERGER^{a}, JILA, University of Colorado and National Institute of Standards and Technology, and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309.

Dihalocarbenes of the form CX_{2} (X = F, Cl, Br, I) are important intermediates in organic reactions, due to the unique chemical properties of their low-lying neutral singlet ^{1}A_{1} and triplet ^{3}B_{1} states that affect their reactivities. Disagreement between experimental and theoretical results for the singlet-triplet splittings in CCl_{2}^{−}, CBr_{2}^{−}, and CI_{2}^{−} prompted a careful reinvestigation of the photoelectron spectra of these compounds, revealing contamination from HCX_{2}^{−} species. Contributions from contaminants previously obscuring the dihalocarbene ^{3}B_{1} triplet spectra have been minimized, and photoelectron spectra of isolated HCX_{2}^{−} compounds were collected to further elucidate the dihalocarbene experimental data. The new CCl_{2}^{−} 351 nm photoelectron spectra exhibit vibrationally resolved transitions to singlet and triplet electronic states; careful analysis drastically alters the previous identification of the triplet state origin, resulting in substantially improved agreement with high level calculations.

aSupport from NSF and AFOSR is gratefully acknowledged